tutor help on this pls pls pls. One number is twice another, and the sum of their reciprocals is 3. What are the numbers?
Answer by Lenzky(18)

Let x be the first number and 2x be the second.
1/x + 1/2x = 3
2/2x + 1/2x = 3
3/2x = 3 multiply both side by 2x
3 = 6x divide by 6 to get the value of x
x = 1/2
CHECKING:
x=1/2 get the reciprocal which is 2
2x= 2(1/2)
2x= 1 get the reciprocal which is 1/1, you can just use 1
then add
2+1= 3